OpenStudy (anonymous):

3a+6a≥6(6-4a)+5(5a-4) help?

OpenStudy (anonymous):

3a+6a>=36-24a+25a-20 9a>=16+a 9a-a>=16 8a>=16 a>=2

OpenStudy (anonymous):

oh thanks! how would I graph it ?

OpenStudy (anonymous):

nvm I know:)

OpenStudy (anonymous):